Calculate the molality of a 100g table sugar (C12H22O11) with a molar mass of 

342.3g/mol placed in 350g of H2O.


1
Expert's answer
2021-03-17T06:45:29-0400

"n(sugar) = \frac{m}{M}= \frac{350}{342.3} = 1.022 \;mol \\

Molality = \frac{moles \;of \;solute}{kg \;of \;solvent} = \frac{1.022 }{0.1}= 10.22 \;mol/kg"